# On a pinning model in correlated Gaussian random environments

### v1

@@ 1.3. Main results
A natural condition for the rescaled Wick ordered partition function to converge weakly is $\alpha+H-1>0$, where $\alpha+H-1$ is the scaling exponent for $\beta$ (see (1.11)). Note that for directed polymers, under the condition that the corresponding scaling exponent is positive, the continuum partition converges in $L^{2}$ as a Wiener chaos (see [6, 7, 19]). However, for our pinning model, assuming $\alpha+H-1>0$ we can only prove that each chaos is a well-defined Skorohod integral which is $L^{2}$-integrable (see Lemma 3.2), and we need a stronger condition $H\in(\frac{1}{2},1),\alpha>\frac{1}{2}$ to guarantee the $L^{2}$-convergence as stated in Theorem 1.1 (also see Remark 3.3 for the criticality of $\alpha=\frac{1}{2}$). This is a new phenomenon that does not appear in the models studied in [1, 6, 18, 19, 7]: in those models, the rescaled partition functions are always uniformly bounded in $L^{2}$ if the models are in the disorder relevant regime, no matter the random environments are independent or correlated. This hints that the pinning model in correlated environment indeed carries some new features.
@@
Motivated by the Weinrib-Halperin prediction, we conjecture that under the condition $\alpha+H-1>0$ with $\alpha\leq\frac{1}{2}$, the rescaled Wick ordered partition function $\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ converges weakly to some $L^{1}$-integrable random variable. We will discuss further on the uniform integrability and the weak limit of $\big{\{}\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big{\}}_{N}$ in Section 3.2.
@@
To close this section, let us highlight our contributions in this paper. Firstly, we obtain some finer property of a disordered pinning model with environment correlation of the form $n^{-p}$, where $p<1$, which complements the result of Berger [2]. To be specific, we find scaling limits of rescaled (Wick ordered) partition functions in the disorder relevant regime. In particular, the Weinrib-Halperin prediction is partially conformed when it differs from the Harris criterion. We also show that the rescaled (Wick ordered) partition functions are not $L^{2}$-bounded in the entire disorder relevance regime (see Remark 3.3), which is a new feature compared to other disordered models. Secondly, we develop a decoupling approach to prove uniform integrability (see Section 3.2), and provide some sharp upper bounds for the norm of traces (see Lemma 4.1) and for the second moments of multiple Stratonovich integrals (see Lemma 4.4). These auxiliary results may be of independent interest.
Notations 1.5.
[…]
@@
The rest of the paper is organized as follows. In Section 2, we introduce some preliminaries. The proofs of Theorem 1.1 and Theorem 1.2 will be presented in Section 3 and Section 4, respectively. We will only deal with the case $10<\alpha<$ 1 in the proofs, noting that the other case $\alpha\geq 1$ can be treated in the same way. Some technical estimates are put in Appendix A.
@@ 2.1. Gaussian spaces
On a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ satisfying usual conditions, let $\dot{W}=\{\dot{W}(t):t\in[0,1]\}$ be a real-valued centered Gaussian noise with covariance
@@
$f,g\right>_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)K(s-t)g(t){\mathrm{d}}s{\mathrm{d}}t=\int_{0}^{1}\int_{0}^{1}f(s)g(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t.$ (2.2)
@@
$\|f\|_{|\mathcal{H}|}^{2}:=\int_{0}^{1}\int_{0}^{1}|f(s)g(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t<\infty.$ (2.3)
@@
$\mathbf{I}_{n}(f^{\otimes n}):=H_{n}(W(f)).$ (2.4)
Let ^${\mathcal{H}}^{\hat{\otimes}k}$ be the subspace of ${\mathcal{H}}^{\otimes k}$ containing only symmetric functions. For $f\in{\mathcal{H}}^{\hat{\otimes}k}$, the $k$-th multiple Wiener integral $\mathbf{I}_{k}(f)$ can be defined by (2.4) via a limiting argument, and moreover,
@@
For general $f\in{\mathcal{H}}^{\otimes k}$, we simply set $\mathbf{I}_{k}(f):=\mathbf{I}_{k}(\hat{f}),$ where $\hat{f}$ denotes the symmetrization of $f$. We also use the following integral form for multiple Wiener integrals:
@@
For square integrable random variables $F$ and $G$ with Wiener chaos expansions $F=\sum_{k=0}^{\infty}\mathbf{I}_{k}(f_{k})$ and $G=\sum_{k=0}^{\infty}\mathbf{I}_{k}(g_{k})$ respectively, we have $F\diamond G=\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\mathbf{I}_{n+m}(f_{n}\otimes g_{m})$, as long as the right-hand side term exists in some proper sense. For a centered Gaussian random variable $F$, its Wick exponential is defined by
@@
The multiple Stratonovich integral of $g^{\otimes n}$ for $g\in\mathcal{H}$ is defined by $\mathbb{I}_{m}(g^{\otimes n}):=W(g)^{n},$ and $\mathbb{I}_{n}(g)$ for $g\in\mathcal{H}^{\hat{\otimes}n}$ by a limiting argument. We also take the following notation for $\mathbb{I}_{n}(f)$
@@
[…]
where $K(x-y)$ is given in (2.1). Multiple Stratonovich integrals $\mathbb{I}_{m}(f)$ and multiple Wiener integrals $\mathbf{I}_{m}(f)$ are related via the following celebrated Hu-Meyer’s formula:
@@
as long as $\vvvert f\vvvert_{n}<\infty$, where
$\vvvert f\vvvert^{2}_{n}:=\mathbb{E}[|\mathbb{I}_{n}(f)|^{2}]=\sum_{k=0}^{[\frac{n}{2}]}\frac{1}{(n-2k)!}\left(\frac{n!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\hat{f}\|^{2}_{\mathcal{H}^{\otimes(n-2k)}}.$ (2.9)
@@ 2.2. U-statistics
[…]
@@
[…]
@@
We introduce the U-statistics for our model, which appears in the chaos expansion for the Wick ordered partition function (1.10).
@@
$\mathbf{I}_{m}^{(N)}(f)$ $:=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right)\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}$ (2.11)
$:=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)(\boldsymbol{t})\,\omega_{\boldsymbol{n}}^{\diamond},$
@@
$\omega_{\boldsymbol{n}}^{\diamond}:=\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}~{}\text{ and }~{}\boldsymbol{t}=(t_{1}\cdots t_{m}):=\frac{\boldsymbol{n}}{N}=\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right).$ (2.12)
@@
$\|\mathcal{A}_{N}(f)\|_{|\mathcal{H}|^{\otimes m}}\leq C^{m}\|f\|_{|\mathcal{H}|^{\otimes m}}$
@@
The desired result follows directly from [19, Lemma C.1]. ∎
@@
For $\boldsymbol{n},\boldsymbol{n}^{\prime}\in\llbracket N\rrbracket^{m}$, by Lemma A.2, we have:
$\mathbb{E}\left[\omega_{\boldsymbol{n}}^{\diamond}~{}\omega_{\boldsymbol{n^{\prime}}}^{\diamond}\right]=\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big{(}n_{i}-n^{\prime}_{\sigma(i)}\big{)},$
where $\mathcal{P}_{m}$ denotes the set of all permutations on $\llbracket m\rrbracket$. Thus, we have
$\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(f)\Big{)}^{2}\Big{]}$ $=\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(\hat{f})\Big{)}^{2}\Big{]}$
$=N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big{(}n_{i}-n^{\prime}_{\sigma(i)}\big{)}$
@@
where the last equality follows from the symmetry of $\mathcal{A}_{N}(\hat{f})$.
@@
By (1.6) and (1.7), we have
$\gamma_{N}(t)\lesssim|t|^{2H-2}\text{ and }\lim\limits_{N\rightarrow\infty}\gamma_{N}(t)=|t|^{2H-2}.$
Therefore, we get by a Riemann sum approximation,
@@
$=m!N^{-2m}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}\gamma_{N}\Big{(}\frac{n_{i}-n^{\prime}_{i}}{N}\Big{)}$
@@
First we consider the case where $f$ is an indicator function: $f(x)=\mathrm{I}_{(s,t]}(x)$ with $s,t\in\Omega_{N}$. Then
@@
$\mathbb{E}_{\omega}\Big{[}\Big{(}\mathbf{I}_{1}^{(N)}(f)\Big{)}^{2}\Big{]}$ $=\mathbb{E}_{\omega}\Big{[}N^{-2H}\Big{(}\sum_{n=1}^{N}\omega_{n}\mathbf{I}_{(Ns,Nt]}(n)\Big{)}^{2}\Big{]}$
$=N^{-2H}\mathbb{E}_{\omega}\Big{[}\sum_{n,n^{\prime}\in(Ns,Nt]}\omega_{n}\omega_{n^{\prime}}\Big{]}$
$=N^{-2H}\sum_{n,n^{\prime}\in(Ns,Nt]}\gamma(n-n^{\prime}).$
@@
$\left(\mathbf{I}_{l_{1}}^{(N)}(f_{1}),\cdots,\mathbf{I}_{l_{k}}^{(N)}(f_{k})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{l_{1}}(f_{1}),\cdots,\mathbf{I}_{l_{k}}(f_{k})\right),\quad\text{as}~{}N\rightarrow\infty.$
@@
$=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1})^{\otimes p_{1}}\otimes\cdots\otimes\mathcal{A}_{N}(h_{l})^{\otimes p_{l}}\omega_{\boldsymbol{n}}^{\diamond}$
$=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}$
By Lemma 2.4, we have $\mathbb{E}[(\mathbf{I}_{1}^{(N)}(h_{i}))^{2}]\leq C\|h_{i}\|_{|\mathcal{H}|}^{2}$ uniformly in $N$. Note that $\mathbf{I}_{1}^{(N)}(h_{i})$ is a centered Gaussian random variable, and thus, we have
@@
Combining the joint convergence in Proposition 2.5, we can apply [19, Lemma B.3] to get

### v2

@@ 1.3. Main results
A natural condition for the rescaled Wick-ordered partition function to converge weakly is $\alpha+H-1>0$, where $\alpha+H-1$ is the scaling exponent for $\beta$ (see (1.12)). Note that for directed polymers, under the condition that the corresponding scaling exponent is positive, the continuum partition converges in $L^{2}$ as a Wiener chaos (see [9, 13, 29]). However, for our pinning model, assuming $\alpha+H-1>0$ we can only prove that each chaos is a well-defined Skorohod integral which is $L^{2}$-integrable (see Lemma 3.2), and we need a stronger condition $H\in(\frac{1}{2},1),\alpha\geq\frac{1}{2}$ to guarantee the $L^{2}$-convergence as stated in Theorem 1.1. On the other hand, when $H\in(\frac{1}{2},1)$, we show in Proposition 3.5 that, assuming $\alpha>2-2H$, $\big\{\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big\}_{N}$ is uniformly integrable. Note that in this situation, when $H\in(\frac{3}{4},1)$, $\alpha$ may be strictly less than $\frac{1}{2}$, and thus $\big\{\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big\}_{N}$ is not $L^{2}$-bounded.
This is a new phenomenon that does not appear in the models studied in [1, 9, 30, 29, 13]: in those models, the rescaled partition functions are always uniformly bounded in $L^{2}$ if the models are in the disorder relevant regime, no matter whether the random environments are independent or correlated. This hints that the pinning model in correlated environment indeed carries some new features.
We note that a phase transition from the $L^{2}$ regime to an $L^{p}$ regime with $1<p<2$ has been studied in the directed polymer model with fixed inverse temperature $\beta$. In particular, [20] showed that, while the polymer measure exhibits the same macroscopic behavior in both the $L^{2}$ and $L^{p}$ regimes, its microscopic behavior differs, as evidenced by distinct exponents in the corresponding local limit theorems. By analogy, we expect that this new phenomenon of the gap between the $L^{2}$ regime and the $L^{1}$ regime in our model should relate to some phase transition.
@@
Motivated by the Weinrib–Halperin prediction and the above discussion, we conjecture that under the condition $\alpha+H-1>0$ with $\alpha\leq\frac{1}{2}$, the rescaled Wick-ordered partition function $\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ converges weakly to some $L^{1}$-integrable random variable for any $\hat{\beta}>0$. We will discuss the conjecture in more detail in Section 3.2.
@@
To close this section, let us highlight our contributions in this paper. Firstly, we obtain some finer property of a disordered pinning model with environment correlation of the form $n^{-p}$, where $p<1$, which complements the result of [4]. To be specific, we find scaling limits of rescaled (Wick-ordered) partition functions in the disorder relevant regime. In particular, the Weinrib–Halperin prediction is partially confirmed when it differs from the Harris criterion. We also show that the rescaled (Wick-ordered) partition functions are not $L^{2}$-bounded in the entire disorder relevance regime (see Remark 3.3), which is a new feature compared to other disordered models. Secondly, we develop a decoupling approach to prove uniform integrability (see Section 3.2), and provide some sharp upper bounds for the norm of traces (see Lemma 4.1) and for the second moments of multiple Stratonovich integrals (see Lemma 4.4). These auxiliary results may be of independent interest.
Notations 1.7.
Let $\mathbb{N}$ denote the set of natural numbers without 0, i.e., $\mathbb{N}:=\{1,2,\cdots\}$ and $\mathbb{N}_{0}:=\{0,1,2,\cdots\}$; for $N\in\mathbb{N}$, $\llbracket N\rrbracket:=\{1,2,\cdots,N\}$; For $\in\mathbb{R}, [a]\left[a\right$]means the greatest integer that is not greater than $a$; $\|\cdot\|$ is used for the Euclidean norm; let $\boldsymbol{k}:=(k_{1},k_{2},\cdots,k_{d}),\boldsymbol{x}:=(x_{1},x_{2},\cdots,x_{d})$ etc. stand for vectors in ${\mathbb{Z}}^{d}$ or $\mathbb{R}^{d}$ depending on the context; we use C to denote a generic positive constant that may vary in different lines; we say $f(x)\lesssim g(x)$ if $f(x)\leq Cg(x)$ for all $x$; we write $f(x)\sim g(x)(as\ x\rightarrow\infty)$ if $\lim_{x\rightarrow\infty}f(x)/g(x)=1$; we use $\overset{(d)}{\longrightarrow}$ to denote the convergence in distribution (also called the weak convergence) for random variables or random vectors. For a random variable X, $\|X\|_{L^{p}}:=(\mathrm{E}\left[|X|^{p}\right])^{1/p}$ for $p\geq 1$. We use $\mathrm{I}_{A}$ to denote the indicator function, i.e., $\mathrm{I}_{A}(\boldsymbol{t})=1$ if $\boldsymbol{t}\in A$ and $\mathrm{I}_{A}(\boldsymbol{t})=0$ if $\boldsymbol{t}\notin A$.
@@
The rest of the paper is organized as follows. In Section 2, we introduce some preliminaries. The proofs of Theorem 1.1 and Theorem 1.3 will be presented in Section 3 and Section 4, respectively. We will only deal with the case $0<\alpha<1$ in the proofs, noting that the other case $\alpha\geq 1$ can be treated in the same way. Some technical estimates are put in Appendix A.
@@ 2.1. Gaussian spaces
On a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ satisfying the usual conditions, let $\dot{W}=\{\dot{W}(t):t\in[0,1]\}$ be a real-valued centered Gaussian noise with covariance
@@
$\left<f,g\right$$_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)K(s-t)g(t){\mathrm{d}}s{\mathrm{d}}t=\int_{0}^{1}\int_{0}^{1}f(s)g(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t.$ (2.2)
@@
$\|f\|_{|\mathcal{H}|}^{2}:=\int_{0}^{1}\int_{0}^{1}|f(s)f(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t<\infty.$ (2.3)
@@
Remark 2.1.
The inner product defined in (2.2) is indeed positive definite since for $H\in(1/2,1)$,
$\left<f,f\right$$_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)f(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t$
$=\int_{0}^{1}\int_{0}^{1}f(s)f(t)\left(C_{H}\int_{\mathbb{R}}|s-r|^{H-\frac{3}{2}}|t-r|^{H-\frac{3}{2}}{\mathrm{d}}r\right){\mathrm{d}}s{\mathrm{d}}t$
$=C_{H}\int_{\mathbb{R}}\left(\int_{0}^{1}f(s)|s-r|^{H-\frac{3}{2}}{\mathrm{d}}s\right)^{2}{\mathrm{d}}r\geq 0.$
Also note that
$\|f\|_{L^{1}(\mathbb{R})}^{2}\leq\int_{0}^{1}\int_{0}^{1}|f(s)f(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t\leq C_{H}\|f\|_{L^{1/H}(\mathbb{R})}^{2},$
where the second inequality is due to Lemma A.4. Hence, we have $L^{1/H}\subset|\mathcal{H}|\subset L^{1}$. For more details, we refer to Section 5.1.3 in [24].
@@
$\mathbf{I}_{n}(f^{\otimes n}):=H_{n}(W(f)),$ (2.4)
where $f^{\otimes n}$ is a function on $0,1]^{n}$ defined by $f^{\otimes n}(t_{1},\cdots,t_{n}):=\prod_{k=1}^{n}f(t_{k})$.
As in (2.2) and (2.3), let the Hilbert space ${\mathcal{H}}^{\otimes k}$ be the completion of smooth functions with compact support under the inner product:
$\left<f,g\right>_{\mathcal{H}^{\otimes k}}$ $=\int_{[0,1]^{2k}}f(s_{1},\cdots,s_{k})g(t_{1},\cdots,t_{k})\prod_{i=1}^{k}|s_{i}-t_{i}|^{2H-2}{\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{t},$ (2.5)
and we also denote by $|\mathcal{H}|^{\otimes k}$ the linear subspace of measurable functions $f$ on $0,1]^{k}$ satisfying
$\|f\|_{|\mathcal{H}|^{\otimes k}}^{2}:=\int_{[0,1]^{2k}}|f(s_{1},\cdots,s_{k})g(t_{1},\cdots,t_{k})|\prod_{i=1}^{k}|s_{i}-t_{i}|^{2H-2}{\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{t}<\infty.$ (2.6)
Let ${\mathcal{H}}^{\hat{\otimes}k}$ be the subspace of ${\mathcal{H}}^{\otimes k}$ containing only symmetric functions. For $f\in{\mathcal{H}}^{\hat{\otimes}k}$, the $k$-th multiple Wiener integral $\mathbf{I}_{k}(f)$ can be defined by (2.4) via a limiting argument, and moreover,
@@
[…]
@@
For square integrable random variables $F$ and $G$ with Wiener chaos expansions $F=\sum_{k=0}^{\infty}\mathbf{I}_{k}(f_{k})$ and $G=\sum_{k=0}^{\infty}\mathbf{I}_{k}(g_{k})$ respectively, we have $F\diamond G=\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\mathbf{I}_{n+m}(f_{n}\otimes g_{m})$, as long as the right-hand side term is well defined (i.e., the series converges in $L^{2}$). For a centered Gaussian random variable $F$, its Wick exponential is defined by
@@
The multiple Stratonovich integral of $g^{\otimes n}$ for $g\in\mathcal{H}$ is defined by $\mathbb{I}_{n}(g^{\otimes n}):=W(g)^{n},$ and $\mathbb{I}_{n}(g)$ for $g\in\mathcal{H}^{\hat{\otimes}n}$ by taking limits for linear combinations (see [18] for more details). We also take the following notation for $\mathbb{I}_{n}(f)$
@@
$\mathrm{Tr}^{k}f(t_{1},.,t_{n-2k})$ $\overset{\text{def}}{=}\int_{[0,1]^{2k}}f(s_{1},.,s_{2k},t_{1},.,t_{n-2k})\prod\limits_{i=1}^{k}K(s_{2i}-s_{2i-1}){\mathrm{d}}s_{2i-1}{\mathrm{d}}s_{2i},$ (2.9)
where K(xy)K(x-y) is given in (2.1). Multiple Stratonovich integrals $\mathbb{I}_{m}(f)$ and multiple Wiener integrals $\mathbf{I}_{m}(f)$ are related via the following celebrated Hu–Meyer formula:
@@
as long as $\interleave f\interleave_{n}<\infty$, where
$\interleave f\interleave^{2}_{n}=\mathbb{E}[|\mathbb{I}_{n}(f)|^{2}]=\sum_{k=0}^{[\frac{n}{2}]}\frac{1}{(n-2k)!}\left(\frac{n!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\hat{f}\|^{2}_{\mathcal{H}^{\otimes(n-2k)}}.$ (2.11)
@@ 2.2. U-statistics
$\mathcal{R}_{N}^{m}:=\Big\{\Big(\frac{\boldsymbol{i}-\vec{1}_{m}}{N},\frac{\boldsymbol{i}}{N}\Big]:\boldsymbol{i}\in\llbracket N\rrbracket^{m},\vec{1}_{m}\text{ is the $m$-dimensional vector of all ones}\Big\}.$
@@
$\mathcal{A}_{N}(f)(\boldsymbol{t})=\sum_{B\in\mathcal{R}_{N}^{m}}\frac{1}{|B|}\int_{B}f(\boldsymbol{s})d\boldsymbol{s}\cdot\mathrm{I}_{B}(\boldsymbol{t}),$ (2.12)
@@
We introduce the U-statistics for our model, which appears in the chaos expansion for the Wick-ordered partition function (1.11).
@@
$\mathbf{I}_{m}^{(N)}(f)$ $=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right)\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}=:N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)(\boldsymbol{t})\,\omega_{\boldsymbol{n}}^{\diamond},$ (2.13)
@@
$\omega_{\boldsymbol{n}}^{\diamond}:=\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}\quad\text{ and }\quad\boldsymbol{t}=(t_{1}\cdots t_{m}):=\frac{\boldsymbol{n}}{N}=\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right).$ (2.14)
@@
$\|\mathcal{A}_{N}(f)\|_{|\mathcal{H}|^{\otimes m}}\leq C^{m}\|f\|_{|\mathcal{H}|^{\otimes m}},$
@@
The desired result follows directly from [29, Lemma C.1]. ∎
@@
For $\boldsymbol{n},\boldsymbol{n}^{\prime}\in\llbracket N\rrbracket^{m}$, by Lemma A.2, we have $\mathbb{E}\left[\omega_{\boldsymbol{n}}^{\diamond}~\omega_{\boldsymbol{n^{\prime}}}^{\diamond}\right]=\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big(n_{i}-n^{\prime}_{\sigma(i)}\big)$, where $\mathcal{P}_{m}$ denotes the set of all permutations on $\llbracket m\rrbracket$. Thus, we have
$\mathbb{E}\Big[\Big(\mathbf{I}_{m}^{(N)}(f)\Big)^{2}\Big]$ $=\mathbb{E}\Big[\Big(\mathbf{I}_{m}^{(N)}(\hat{f})\Big)^{2}\Big]=N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big(n_{i}-n^{\prime}_{\sigma(i)}\big)$
@@
[…]
@@
By (1.6) and (1.7), we have $\gamma_{N}\lesssim|t|^{2H-2}$ and $\lim_{N\rightarrow\infty}\gamma_{N}=|t|^{2H-2}$. Therefore, we get by a Riemann sum approximation,
@@
First we consider indicator functions of the form $f(x)=\mathrm{I}_{(s,t]}(x)$ with $s,t\in\Omega_{N}$. Then
@@
$\mathbb{E}_{\omega}\Big[\Big(\mathbf{I}_{1}^{(N)}(f)\Big)^{2}\Big]$ $=\mathbb{E}_{\omega}\Big[N^{-2H}\Big(\sum_{n=1}^{N}\omega_{n}\mathbf{I}_{(Ns,Nt]}(n)\Big)^{2}\Big]=N^{-2H}\sum_{n,n^{\prime}\in(Ns,Nt]}\gamma(n-n^{\prime}).$
@@
$\left(\mathbf{I}_{l_{1}}^{(N)}(f_{1}),\cdots,\mathbf{I}_{l_{k}}^{(N)}(f_{k})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{l_{1}}(f_{1}),\cdots,\mathbf{I}_{l_{k}}(f_{k})\right),\quad\text{as}~N\rightarrow\infty.$ (2.17)
@@
$=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1})^{\otimes p_{1}}\otimes\cdots\otimes\mathcal{A}_{N}(h_{l})^{\otimes p_{l}}\omega_{\boldsymbol{n}}^{\diamond}=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}$
By Lemma 2.5, we have $\mathbb{E}[(\mathbf{I}_{1}^{(N)}(h_{i}))^{2}]\leq C\|h_{i}\|_{|\mathcal{H}|}^{2}$ uniformly in $N$. Note that $\mathbf{I}_{1}^{(N)}(h_{i})$ is a centered Gaussian random variable, and thus its higher moments can be expressed via its variance. Hence, we have for any $q\in\mathbb{N}$,
@@
Combining the joint convergence in Proposition 2.6, we can apply [29, Lemma B.3] to get